[8] Plot the following complex number in the complex plane, write it in long-hand polar form with the argument in degrees,
Answer #1

Q8 4 5 cio 13T 18 Long hand 5 cos 13 T + i 5 sin 137 18 18 + 3.8302222 rectingular - 3.213938 (0= 137, 8=5) r 8-5 (2 J3 +2134x6 = 641 i (2 KTT + 71/2) 26 216 K=0,1,2,3,4,5 a 2 ei (KI + I) zei (4 kI + 12 x = a e 26=2 cis (7) H, = 2015 (51) (977), 237